Day shift: 9am–9pm /  Night shift : 9pm–9am

ThursdayNov/1Dismantling previous experiment. ESA setup preparation starts 10am. 
FridayNov/2Day: preparation  /  Night: first beam ~9pm 
SaturdayNov/3Day: beam run  / Night:  MD, no beam to ESTBInitial test runs
SundayNov/4Day: ESTB beam / Night: no ESTB beamComplete setup / first runs
MondayNov/5Day: ESTB beam / Night: no ESTB beam2x LBNL devices at 0 degrees
TuesdayNov/6Day: ESTB beam  / Night: 2 extra hours of beam  (9AM-11PM)2x LBNL devices at inclined
WednesdayNov/7Day: Late start ESTB beam at 11AM / Night: no ESTB beam (11AM-9PM)2x SLAC devices

Main goals:

  • Test 4 RD53A single chip modules with different types of planar sensors from LBNL and SLAC.
  • Characterize spatial resolution for difference pixel sizes 50x50 and 25x100, different active thickness 100um and 150um.
  • Absolute charge calibration for MIP hits
  • Diagnose L1 time dispersion
  • Further studies of different RD53A FE types and efficiencies,   
  • Commission LBNL cosmic telescope in beam 

Preparation items:

1) Support table/platform behind Caladium to prop up the LBNL cosmic telescope of 1x1 foot cross section 3 feet long (with ~2 feet height adjustment).

2) Wireless boost in the tunnel:  Juan from TestFac has been working on this and will ensure its installation.

3) Beam trigger availability: Will have 2-3 independently controllable beam triggers to work with (TTL or NIM). Will also add a NIM fanout in case needed.     .  

4) We ran short of remote CAM last time. Some more CAMs resurrected ?

5) Need to refresh/document procedure to operate the remote control of PSU->HV. Adding a couple of channels ?


Schedule discussion with Carsten:  Thursday 9am beam off to dismantle previous expt. SD will meet Carsten 10am at ESA to start process of moving Caladium back on beam line and decide how long a section of the beam pipe should be taken out (next experiment will need them reinstalled). Will also look for the support platform for LBNL cosmic telescope. Also confirmed we need the Wednesday day shift beam at the end.  


Our plan declared  at 4:15pm weekly planning meeting. Learnt that the Sunday-Tuesday night shift MEC experiment LU57 is out of the usual ESTB friendly mode for this period with special ultra-stable beam requirement which will take the full 120Hz and not be able to split off 5Hz for ESTB. Friday night shift with XPP is still OK for ESTB, while Saturday-Wednesday Day shifts will have the regular 5Hz 11 GeV beam for ESTB. Friday-Saturday will run with whatever ESTB setup known previously, while Sunday onwards can get dedicated beam tuning from Toni Smith.


Preparation started at ESA:

  • A remote controllable XY stage lifted into ESTB tunnel behind Caladium. This is intended to support the LBNL cosmic telescope in horizontal position (1x1 foot cross section 4 feet long) and followed by another silicon sensor Carsten wanted to mount right behind it. This needs a total of ~56 inch length between the two objects. Carl found a metal plate to stack on the top deck to extend the platform longitudinally. The top surface is ~20cm below the beamline. The LBNL telescope can probably lie on the deck with the feet retracted fully up.
  • Doug checked the XY drives of the back stage (Beckoffs BKH21: B061 Big Bertha) and they seem to be functional. Y move very slow. Not sure where we are on range. 
  • Juan installed wireless router in the tunnel at middle entrance - visitor network strength full bar.
  • Placed 7 web cams around Caladium. The AXIS types are back to life again (we missed the power supplies previously).
  • One HSIO2 already in place for DUT readout when needed. 


Plan of the day: 


  • Move Caladium into beamline and try dummy runs. 
  • Drill holes on top plate of the back stage and prepare L anchor on Bosch frame to get ready for mounting the LBNL telescope
  • Mount Carsten's silicon sensor on mini stage onto the back stage   
  • Fix focus of two AXIS web cams (13,14). 
  • Check EPICS remote control of PSUs and possibly add some channels.
  • Exercise backstage XY drive to map out the adjustable range
  • Locate all 3 unused DUT DAQ Ethernet lines and test+tag them 
  • Locate additional beam trigger cable(s) and install NIM fanout and NIM-TTL in case more separate triggers needed  


  • LBNL party arrives ~1pm after fueled with dim sum
  • 2-3pm: half of LBNL party in 116PRA. Others+SLAC continue installation
  • 3:45pm ESTB orientation (AD106 training) for all new comers
  • Dry runs without beam


Caladium beamline move postponed until first beam run through. Debated whether that needs to be primary beam or not. Eventually MCC will just try secondary to avoid the need of RP survey. 

Dummy runs with Caladium only worked out smoothly ex expected.

Carsten's (actually Chris Kenney's sensor) silicon sensor will be mounted near the dump area instead and not interfering with our setup.

LBNL party arrived 10min before the controlled access search so just moved some equipment into ESA and then exit to allow the search.

ESTB orientation AD106 done by Keith for 6 new people.

RD53A test setup moved along but not completely done yet until tomorrow. The beam trigger is Teed to fanout several copy of it for the L1 timing tests so that TLU entrance signal is slightly delayed than usual. Cosmic telescope modules are mounted but powering/readout not yet connected. One LV banana plug looks like is missing the tightening screw. 

The backstage X motor exercised and looked like we can move all the way between the two stoppers with at least 50cm worth of travel range. We moved the stage to be ~6-7cm from the wall side stopper so that the vertical post on the stage is also on the wall side with 5cm clearance from beam. The stage drive can move +x to bring the cosmic telescope back to beam line when the telescope feet are placed on the marked positions.   

~10pm: exited ESA to let MCC try secondary beam.

10:30pm MCC thought they are close to be done with tuning so we started to move Caladium coarse X in reduced x direction and running the Caladium only readout. Before reaching the beam spot, the secondary scatter was ~50 particles/shot but once beam spot came into the frame, that turned to be more like 2000/crossing. Beam spot was also very wide in X ( more than the 2cm MIMOSA aperture). Asked MCC to reduce rate and spot size. a few tries didn't make much impression until a move to close down SL-10 to -0.5. That brought the rate down to ~100 particles/crossing. Beam spot is filling good fraction of aperture ans looks like a good tuning setup to start with tomorrow. The Caladium XYstage X/Ycoarse RBV=115.4/3.8mm (Y is very close to min limit). Run 336 tuning beam spot picture below. 

Done for tonight      


Plan of the day: 

9am-9pm Day shift will continue with 5Hz beam running with XPP as primary expt. We expect to start ~9am for access to continue towards full setup for RD53A and cosmic telescope then run until 9pm.  

Activities:  Timon, Ben, Su Dong, Zijun, Ke, Sanha 

9:30: Access to continue setup started.

10:30  FE-I4 telescope installation saw some sensitive HV current fluctuations on on module. The module box not light tight enough ? Some Al foil wrap shield away light helped somewhat. Looked fro CAT6 cables at ESA and SLAC to instrument FE-I4 data/Cmd connections. Cables are of somewhat different types. Last plane (5) had trouble with digital calib initially but worked after replacing with another data cable.     

11:30  Moved the back XY-stage to +ve X into the beam line. Also adjusted the Y move down by 3-4cm into the beam line.  The horizontal laser battery ran out which we replaced but the push of it made it very difficult to reinstall sot hat it is temporarily mounted on the frame next to beam exit window and its tilt changes with rotation so cannot be relied on for the moment.

12:30  Completed installation. Exited ESA and requested beam.

12:45  Operator reactivated the same setup from last night. Beam spot still at the same place and size as last night, but particle rate is down to ~50.  Will keep going with this.

12:45  Moved back stage X by -75mm to take the FEI4 modules out of beamline for calibration and threshold tuning.

13:00  Carsten called. Lead Glass counters (Scope ESA brown trace) was left at 900V and 10mV/div scope display last night in anticipation of primary beam. He raised the HV (to 2K?) and scope back 50mv/div.  Signal size 50-100 mV for the ~50 particles per shot. He moved the counter slightly to verify signal went away so it looked like crudely beam doesn't have diffused tail.    

14:00 Full set of 3000e threshold tunes obtained for the LBNL cosmic telescope modules and they are temporarily stored at /home/beam/daq/rce/moduleconfigs/fei4bexample/ directory on esadutdaq1 for the RCE DAQ calibrations with the 6 modules give a temp  name 94-22-0x (x=0-5) until their real IDs are known.   

15:00  Created the EUDET configuration LBNL_cosmic_telescope.conf to include all 6 planes of cosmic telescope as DUT. After a few errors we learned that FEID, SensorID also need to have distinctive values to make EUDAQ really happy.  

19:00  A long access to figure out RD53A setup. Early setup for 0x0A57 (cable D) stopped working while 0x0A53 (cable C) always had trouble. This looked like an issue with the new PCIe card for 0x0A53 might not have the fully compatible firmware. Recovered D while tomorrow is schedule to debug out C.    

19:30  Added a LEMO between the first module of cosmic telescope HitOR and HSIO2 LEMO bank chan #1 (2nd socket from back edge) hoping to run the telescope in HitOR triggered mode. This is not strictly correct electrically as HitOR is 1.2V LVCMOS while HSIO2 LEMO port expects TTL. Trying our luck here.   

Select data worth looking at:

~3:00 PM Run 346. A long run with the LBNL cosmic telescope and full MIMOSA. Su Dong, Zijun, and Sanha found out from this run that the second FEI4B module of the LBNL cosmic telescope wasn't outputting anything. This module was almost completely masked out during the fasttune process. Sanha tried re-calibrating and re-tuning the module for a while, but could not fix it. Zijun spotted a lose HV connector on this module but that didn't solve the problem. Data cable swapped with the 3rd module (block CAT6e) still didn't make much impression. More attempts to calibrate didn't get around the problem such that bypassing the crosstalk stage only delay the masking out to the noise test stage. Another issues was that many modules got a funny vertical strips of bad pixels on the top right corner from calibration in a similar manner. Timon asked us to check possible RCE readout errors, but this might be more likely due to the possible tangle of the 6 modules sharing the same fei4bexample 'module' folder ?      

~3:50 PM Run 350. Timon got a first successful run with one RD53A module. This was with device 0x0A57 on cable D and cable B to TLU DUT #3 (mask +0x8).  

from 7:45 PM to 8:25 PM Run 400. A long run with one RD53A module (0x0A57) and full MIMOSA (full 10k events). The trigger latency parameter for YARR adjusted slightly to put the L1ID closer to the center of DAQ buffer. The beam did not cover the entire RD53A sensor area, so Ke and Sanha tried tuning the fine position to move the module around.

from 8:50 PM to 9:00 PM Run 407. A short run (due to beam shutdown) with somewhat better aligned RD53A position. The short-side (horizontal direction) is well-covered by the beam. On the long-side (vertical), the beam still seems to be 100~150 pixels off the center; the module has to be moved 5~7.5mm down (up...?). Note that the beam shape is extremely asymmetric in the vertical-direction. Current RBV config: X coarse 115.41mm, Y coarse 3.83mm, Y fine 7.69mm, X fine -56.53mm, Rotation 213.51mm

21:00  Beam off 


9am-9pm beam for ESTB

Plan of the day: 

  • Debug the RD53A 0x0A53 (C) setup 
  • Special timing tests 
  • Cosmic telescope self-trigger mode ? Recovering 2nd module ? 
  • Systematic RD53a data taking runs 
  • Cosmic telescope data runs to map out all device areas   

Activities:  Ben, Timon, Adrien, Patrick, Sanha, SD

9:45 AM Beam turned on with ~40 hits per beam on MIMOSA

9:45 AM to 10:00 AM Run 408. Su Dong and Sanha tuned the telescope and the RD53A location to center the beam. The bottom ~30% (x=0-108 Synchronous FE causing noise to other FEs) of the RD53A is not configured for read-out, so we focus on covering 130px and above only (center the beam at mean roughly 260px).

10:00 AM to 10:20 AM Run 409. Su Dong and Sanha running a long run with the tuned telescope & RD53A location. 40 hits per beam on MIMOSA, 15 hits per beam on RD53A. This rate seems too high, so Su Dong talked to the beam control team to reduce the beam density, while keeping the same coverage.

10:25 AM  Discussion with Ben decided that we should aim for 1 particle in RD53A per shot. We are 2-3 times high wrt that. The beam spot is nicely spread over almost the full MiMOSA area quite uniformly so that we should try to preserve. Called MCC to request reducing rate which they offered to try inserting the screens which didn't seem to make that much difference until the last screen. Carsten called to complain the strategy as he believes inserting screen after SL-10 might distort the exit particle moment spectrum to let in some potentially unwanted low momentum particles. Carsten talked to MCC for a revised configuration to work on the SL-10 slit to narrow the momentum phase space but hopefully not disturbing the large uniform spot. This appeared to have returned us to the previous state with 8-9 particles on MIMOSA. Called MCC to try narrow SL-10 from -0.6mm to 0.7mm. This lowered the MIMOSA rate to 2 or 3 particles while RD53a was ~0.6. So we tried to go back some to 0.65mm which brought the rate back up to (cluster multiplicity) of ~6 and RD53a having ~1.5. Decided to keep running with this. Carsten noted that the SL10 readback was still front/back 0.69,0.7 which implies that the SL10 drive to 0.65 is beyond the actual drive precision (SL10 is a very heavy object) not nudged it enough to make the difference. For further tweaking of the rate up, Carsten recommended only driving only the back edge of SL10.               

11:00  Run 418 with the new beam setup started EUDET+RD53A

11:35  Run 420 started with same conditions

12:50  Runs 420 and 423 completed (10k events each) with the same condition

14:00  Run 426 reached 10K and EUDAQ tried to auto start another run but that process crashed YARRproducer. Restarted Run control and new run 428. 

14:20  Calibration for FEI4 telescope progressed with module 2 recovered, but some other modules showing strange results. Noise scans with 4 modules crashes calibserver. This was eventually overcome and a full set of calibrations of all 5 modules stored in their own directories modconfigs/lbnl_cosmic/mod* (*= 0-5). cfg #27  was the eventual state reached that should be used for default running.     

14:25  Ended Run 428 for access. RD53A timing setup adjustments, Added a Lemo-T to the TLU beam trigger input to send a signal to HSIO2 Ext trigger LEMO port to allow cosmic telescope to be driven by the beam trigger independently with cosmicGui. Also took out the lemo on first module's HitOR just in case that might be a noise source.

15:10  Restarted datataking after access with same setup as before with single chip for Run 430. 

15:30  Ended Run 430 to add 2nd RD53a.  

15:35  Started Run 434 with both RD53a in DUT and data looked good. The L1 time between the two devices differed by a few BX, not clear why.  

16:25  MCC called for a short beam off which we just ran through.

16:26: Started Run 441 with both RD53a with two FE-ID set distinctively. Offline checks confirmed hits position correlation between RD53a and MIMOSA.

16:50  Tried to run Cosmic Telescope with cosmicGui with "scintillator trigger" being the beam trigger over LEMO, but no triggers seen. Upon checking Event Log from Sep/2016 session, it was noted there that the beam trigger came over with BNC cable was NIM. HSIO2 LEMO trigger is expecting TTL. 

17:56  End Run 447 to request for access. Changed the RD53a scope monitor scale and brought the LEMO T of beam trigger to an NIM->TTL module and led the TTL LEMO to HSIO-2 TRIG1 port. Also replaced the data/cmd cable for cosmic telescope module 2 which has somewhat ragged noise.        

18:20  Upon restart, the beam trigger no longer was no longer firing. Had to access again to back out the beam trigger T to NIM-TTL for the cosmic telescope. The 2016/Sep Even log did say the NIM signal on the BNC cable for the TLU was already rather marginal in amplitude and the was the last feather broke the camel. We probably should undo the T spaghetti on the beam trigger to use the NIM fanout to distribute replicated NIM signals cleanly.    

19:18  Started Run 456 with beam trigger working again. We followed the written procedure of start_1 and found that only got one RD53a (SPEC#1) going. Texted Timon that we need to run start_0 in addition to bring up SPEC#0. 

19:25  Ended 456 to start Run 457 with both RD53a.  

17:55  Tried to run cosmicGui but couldn't even get the cyclic triggers to run. Looking at the date of the cosmicGui image, it was from Sep/10 when Matthias first ported a working rceProducer with EUDAQ-1.7, but the CosmiGui code with EUDAQ dependence was surely not updated yet, as we knew the version running in the lab needed a major overhaul for this a few weeks ago. This has to be checked with Martin/Matthias to revisit tomorrow.   

20:30  Checked the calibration for cosmic telescope module 2 (since cable was changed). This is should a more uniform noise and the jagged shoulder disappeared. Calbration _56 looks like a good default for module 2. 

20:34  Run 458 ended after reaching 10K events and auto started Run 459.

21:05  MCC called to take away beam. Ended Run 459.  Done for the day.

Summary of good runs. Particles rates estimates ~ clusters/events for MIMOSA/RD53a

RunStartEnd time#eventsRD53a DUTParticle rateComments
408 10:02 D25 / 10Rate slightly high at MIMOSA ~25 particles/xing
409 10:20 D25 / 10 
418 11:3410KD 6 / 1.5After requesting MCC to reduce rate
420 12:0910KD 6 / 1.5 
423 12:4810KD 6 / 1.5 
425 13:2810KD 6 / 1.5 




 6 / 1.5 
428 14:2610KD 6 / 1.5 
430 15:309KD 5 / 1.2Just after access
434 15:372KC,D 5 / 1.5 RD53a_1 (C) has ~10 clus/evt
435 15:59 C,D  
441 16:35 C,D RD53a module ID more distinctive
445 17:07 C,D  
446 17:4110K


447 17:585KC,D5 / 1.53? RD53a DUT mon folders RD53a_200 always empty
456 19:242KD Missed Yarrproducer0 at restart
457 19:5910KC,D  
458 20:3310KC,D6 / 1.9 
459 21.04


C,D6.6 / 2.1 



Plan of the day

  • Continued accumulating statistics with the current 2 LBNL RD53a modules at 0 degrees
  • Investigate cosmicGui setup and routing beam trigger to HSIO2 for independent Cosmic telescope runs  
  • Continue developing the beam vs 40Mhz clock phasing measurement  

Activities:   Ben, Adrien, Elisabetta, Cesar, Sanha, SD 

9:21  Started Run 460 with both RD53a. Same rate as yesterday.

9:56  Ended 460 with 10K events. Started 461.

10:30 Ended 461 with 10K  and Run 462 was autostarted but there was an event mismatch that killed all connections but the monitor and log collector. Run was terminated and restarted.

10:32 After run control was restarted, started Run 463.

11:06 Ended 463 with 10K events, Run 464 autostarted with no issues.

11:38 Ended 464 a teensy early with 9.5K events at Ben's advice in order to avoid event mismatch errors.Will continue to end runs a bit before 10K to avoid these errors that appear because of event mismatches during autostart of runs.

11:39 Started Run 465.

12:12 Stopped 465 and started Run 466.

12:46  Ended 466 with 10K and accidentally ran into autostart, procedure for stopping before 10K to avoid errors will resume with 467.

12:47 Run 467 autostarted.

13:18 Stopped Run 467. Started Run 468.

13:52 Stopped Run 468. Started Run 469.

14:25 Stopped Run 469. Started Run 470.

14:58 Stopped Run 470.

15:01 Started Run 471.

15:34 Stopped Run 471. Requested access from MCC to fix the trigger for the cosmic telescope. Got cosmicGui running with cyclic triggers after realizing that the key was to adapting a working version of .cosmicDaq.rc. Got clarification from Martin on the HSIO2 cosmic trigger connection scheme: a) 4 LEMO ports on RTM are NIM; b) main board LEMO Ext 1,2 are TTL; c) main board Ext 7,8 take LVTTL (2.5V) are Ored with 1,2; d) TLU RJ45 takes TLU triggers. Found trigger 11 (NIM) used previously by UCSC and connected it to the outermost RTM scintillator trigger port on the HSIO2 board (scintillator0). Also connected LEMO from TLU TTL out Trigger 1 to HSIO2 main board Ext Trg 1. This needs to activate in EUDAQ config to run parasitic to main partition when the is up. Don't have level shifter in hand to adapt the HitOR trigger from cosmic telescope for now.    

16:03 Started Run 472 after access.

16:03 cosmicRun 034 First data in LBNL Telescope via cosmicGui. After coming back upstairs and changing the trigger settings in the cosmicGui, we were able to read out data from the cosmic telescope (see cosmicRun 034). The ESTB beam trigger delay (UCSC LGAD trigger) is set at -104500, and the trigger delay on cosmicGui is set at 228, and telescope/DUT latency set to 10 (reversed def for FEI4, low number means high latency).  

The timing distributions of the hits in cosmicRun 034 peaks at two different bins: a small peak at bin1 and a large peak at bin5. Decreasing the cosmic trigger delay by 4 bins (to 224), we see that both the peaks move by 4 bins, suggesting that the two-peak feature is not due to the detector noise (see cosmicRun 035).  Increasing the cosmicGui trigger deadtime to a large value (100 ticks) to avoid double-triggers, we still see essentially the same behavior in cosmicRun 036. After a longer run (cosmicRun 037) and some analyses & discussions, we infer that the beam is indeed split into two bunches, roughly 100ns apart. We also suspect that certain "bad" pixels might be reading out multiple times after it gets hit by real particles (roughly every 100ps); these pixels have very high hits in only the "second bunch" of the beam (but not the first), and only these pixels read out hits well after the beam is gone (well after the second bunch peak). The behavior seems to happen across all the 6 modules.

16:36 Ended Run 472.

16:37 Started Run 473.

17:10 Ended Run 473.

17:11 Started Run 474.

17:44 Ended Run 474.

17:45 Started Run 475.

18:01 Started cosmicRun 038 with beam trigger delay at -104500 and cosmicGui trigger delay at 224.

18:18 Ended Run 475.

18:19 Started Run 476.

18:40 Data taking stopped abruptly for both the cosmic telescope and the two RD53A modules. Couldn't ssh into computer downstairs that starts the TLU Producer. Shut off beam and requested access to see what happened to computer.

18:45 Su Dong and Sanha accessed the hall to power-cycle the computers downstairs. Suspecting that this could have been a heat issue, we put a large fan blowing at the esadutdaq1 computer.

19:20 ar-eudaq machine was also frozen, so it was reboot.

19:45 Successful data taking resumed for RD53A

20:17 Started cosmicRun 051. LBNL cosmic telescope moved around slightly to better center the beam: current config X stage -3.50mm and Y-stage -83.99mm. cosmicGui trigger delay is also set to 216 now, in order to capture just the first bunch of the beam and avoid the suspected "double-ticking" effects by some bad pixels.

21:01 Ended cosmicRun 051 due to the end of beam time.


Data summary:

RunStartEnd time#eventsRD53a DUTParticle rateComments


C,D4.6 / 1.1Start Day shift with beam setup as yesterday
4619:5610:3010KC,D After this run all connections but the monitor and log collector died and the next run was not autostarted. Saved data from run 461, terminated processes, and restarted.
46310:3211:0610KC,D -
46411:0611:389.5KC,D Stopped a little early to avoid potential event mismatch errors, will continue to do this for the rest of the runs.
46511:3912:129.8KC,D -
46612:1212:4610KC,D Accidentally ran to 10K and autostarted next run (with no issues) because we were looking at the cosmicGui and lost track of the count. Stopping a bit before 10K resumes with 467.
46712:4713:189.5KC,D -
46813:1913:519.5KC,D -
46913:5214:249.8KC,D3.9 / 0.68


47014:2514:589.8KC,D -
47115:0115:349.7KC,D -
47216:0316:3610KC,D Started Run 472 after access to ESA to establish trigger for cosmic telescope, decided to let the run control autostart runs again. 
47316:3717:1010KC,D -
47417:1117:4410KC,D -
47517:4518:1810KC,D -
47618:1918:403KC,D Data taking stopped abruptly for both the cosmic telescope and the two RD53A modules. Couldn't ssh into computer downstairs that starts the TLU Producer. Shut off beam and requested access.
47819:4520:2110KC, D Another full 10K run with the same beam setting after accessing due to computer shutdown issues

C, D

48020:559:011668C, D Run ended due to the end of beam time

Nov 6

Note:  ESTB period is  9:00am–11:00pm  today with extra 2 hours   (Wednesday will be shorter 11am–9pm) 

Plan of the Day:                 

  • Switch to tilted data taking with the two LBNL RD53a
  • More cosmic telescope data with separate cosmicGui partition 
  • DAQ clock phase vs beam tests 
  • Depletion HV scan 

Activities   (Adrien, Timon, Ben, Maurice, Charilou, Patrick, Mark, Sanha, SD) 

9:30 Sanha called the beam control room to start the beam. There is a water leakage in the klystrons, and the repair/replacement will take 3-4 hours. Test beam data taking will be postponed until the repair/replacement is complete (probably in the early afternoon).

9:50 cosmicRun 051. Testing the current telescope configuration for any random noises. Cyclic trigger set to 100Hz. Almost no random noise is seen after 100s.

11:00 RD53a DUT stage slight rotated to get slight charge sharing in "phi". Tape measured distances derived an angle of 11.8o. Reconstructed angle from photo gave 13o  which are consistent with errors.  


RD53a DUT position before Nov/6 (perpendicular)RD53a DUT position before Nov/6 (zoom)Nov/6 DUT after rotation

11:50 Beam is back on (earlier than expected!)

11:52 Started data taking

12:35 Channel 10 beam trigger disappeared while Ch 11 was driving the cosmic telescope at 5 Hz happily. Tried to run ch 10 trigger with a generated 10Hz from the trigger module and that also worked. Call Carsten and Mike Dunning and both had no idea how that could happen. The trigger unit was clearly receiving the accelerator signal correctly to run ch 11 while the down stream part delivering trigger to ESTB was also fine with the generated 10Hz.   

13:05 When in for access to see what's going on. Swapping the trigger channels 10 (originally feeding the EUDET) and 11 (originally feeding the cosmic telescope). The original trigger settings are: width 200ns, delay 110488ns for channel 10 (EUDET) and width 100ns, delay 104500ns for channel 11 (cosmic telescope). This didn't work and quickly found out that Ch 10 is NIM but Ch 11 is TTL so we cannot really swap. So we are back to the original pattern of ch 10 running TLU. The Ch 10 NIM pulse looked a bit marginal with slow ramps barely getting to asym level. Inceasred the width to 300ns and also noticed some ripple after pulsing. Cleanup up the setup so that CH 10 BNC swicthed to LEMO to go in a NIM Fanout unit  input and much more polished square pulse output from the Fanout is sent to the TLU. The fanout unit Thr is tweaked to wipe out the after ripple pulses. This cleanup seemed to have worked, but not clear where things fell exactly before this.       

14:55 Beam back after some struggle at MCC with turning on the A-line magnet. Started Run 497.    

15:29 Noticed the beam was off but trigger was still firing because magnet was still on, not sure for how long but there will be less data in Run 497 because although the number of events kept increasing there were no hits on RD53a bc the beam was off.

15:31 Full-ish(10K but not really) run 497 ended and Run 498 started.

16:05 Ended run 498 with full 10K events and started Run 499.

16:39 Ended run 499 with full 10K events and autostarted Run 500 but got an event mismatch error that killed all of the producers.

16:42 Started Run 501.

17:15 Ended run 501 with full 10K events and started Run 502.

17:34 Ended Run 502 with ~4K events.

17:36 Called MCC to increase rate of electrons from 6 per pulse to 600 per pulse.

17:44 Started Run 505 and will try to run Picoscope during this run. Also since the rate is increased, the runs will only be taken until 2K.

17:51 Stopped Run 505 with 2K hits. Picoscope worked well and will continue to use it. Beam went down momentarily and came back up in a couple of minutes.

17:56 Beam came up and started Picoscope readout before starting data taking for Run 506. Will continue to stop runs at 2K events, Procedure for using picoscope after a run: kill the picoscope program, take down the last 5 digits of the picoscope data file corresponding to the run (ll ~/slactb_scope and take the most recent files last 5 digits) and then start the picoscope program again(build/bin/slactb_scope) and start a new run in RC panel.

19:09 Stopped runs to try and get synchronous front end working.

19:59 Attempts to fix the synchronous front end failed, after returning settings to how it was before trying to fix the front end the chip won't turn on. After many attempts to remotely power cycle and fix the problem, an access was requested to go down and fix it manually.

22:11 After graphing the ToT from runs 522-530, noticed that it was not changing as much as expected, leading us to believe that maybe remotely changing the HV might not be working. To check, we set the voltage to 0 and took data but did not see what we expected. The camera pointed at the equipment rack didn't have the HV PS in the frame so we requested an access to go down and see whether remotely chaning the HV actually works or not. We confirmed that the HV was changing but the script must have an indexing error on its for loop because whenever the HV is set, it always ends up at a value 1 off from what was set. For example if we were at 20V and set it to 15 we would end at 16 and if we were at 10 and set to 15 we would end up at 14. There is also an Errcode flashing n Keithley at each stage. After figuring out that the voltage was changing but not to where we expected, a new webcam was moved to have the HV panel in view so it could be monitored.

22:52 Beam restored and Run 534 started.

23:03 MC called to take away beam after Run 534 was completed.


RunStartEnd time#eventsRD53a DUTParticle rateComments
48111:5211:561kC, D RD 53A tilted at 11.8 degrees. Run ended early to change YARR config.
48211:5612:357kC, D4.8 / 0.8The trigger from the machine stopped. Run ended.
49714:5515:3110kC, D  
49815:3116:0510kC, D  
49916:0516:3910kC, D  
50116:4217:1510kC, D  
50217:1517:344kC, D  
50517:4417:512kC, D Rate was increased from 6 e- to 600e- per pulse.
50617:5618:032kC,D Picoscope data file: 55765
50718:0418:122kC,D Picoscope data file: 56275
50818:1218:192kC,D Picoscope data file: 56748  Cluster ToT for 200/210: 7.40/10.02
50918:3018:372kC,D Picoscope data file: 57782
51018:3918:462kC,D Picoscope data file: 58361
51118:4718:542kC,D Picoscope data file: 58824
51218:5519:022kC,D Picoscope data file: 59300
51319:0219:092kC,D Picoscope data file: 59750
52220:3320:402kC,D350&270(hits/plane)Set HV to 61. Cluster ToT for 200/210: 7.29/10.03

Set HV to 41. Cluster ToT for 200/210: 7.36/10.02


Set HV to 31. Cluster ToT for 200/210: 7.50/10.01

52621:0121:082kC,D340&280(hits/plane)Set HV to 21. Cluster ToT for 200/210: 7.45/10.01
52721:1121:182kC,D340&280(hits/plane)Set HV to 15. Cluster ToT for 200/210: 7.53/9.97
52821:1921:262kC,D340&280(hits/plane)Set HV to 11. Cluster ToT for 200/210: 7.49/9.98
52921:2821:352kC,D340&290(hits/plane)Set HV to 7. Cluster ToT for 200/210: 7.41/9.83
53021:3621:432kC,D340&310(hits/plane)Set HV to 3. Cluster ToT for 200/210: 7.36/9.66
53121:5021:572kC,D350&270(hits/plane)Set HV to 51. Cluster ToT for 200/210: 7.37/10.00
53221:5922:062kC,D350&280(hits/plane)Set HV to 1. Cluster ToT for 200/210: 7.50/10.08
53422:5222:592kC,D300&220(hits/plane)Set HV to 0. Cluster ToT for 200/210: 6.08/7.04
Cosmic RunStartEnd time#eventsTrigger DelayComments
051~9:50~9:511k-104500, 224Testing without the beam for random noises. No significant noise seen.
05311:56 12:2710k-104500, 224 
05412:34 13:059k-Cosmic run ended in order to change the trigger over to the RD 53A's
05714:5815:00~600-Short run to make sure trigger/timing is still working as expected (Runs 55 and 56 before this were empty due to beam being off for access to fix triggers)
062   - 



ESTB beam **11am–9pm**     

Plan of the Day

  • SLAC Rd53a device 0x0A97 (25x100 pixel 150um active thickness) replaces the current channel D 0x0A57 (50x50) early enough in the morning to allow Timon tuning it remotely 
  • Setup a cleaned up TTL beam trigger from Ch-11 for cosmic telescope to see if the double timing peaks still exist.
  • Main data taking runs will be steep tilted 25x100 sensors. One DUT at a time, first 0x0A54 (100um LBNL) then 0x0A97 (150um SLAC). 

Activities:  Adrien, Mark, Charilou, Cesar, SD 

10:30 Replaced 0x0A58(labelled A57 in the configs) with SLAC RD53A device 0x0A97.

11:12 Beam was up and started taking runs, but no events were being built, called Timon and he said module A54 does not like to be power cycled. So we requested access to go down and fix it. Fortunately on power cycle was sufficient to bring A54 back. 

11:47 Beam was turned back on after module started working again and Run 541 was started with just one Yarr producer in the run control for the LBL RD53A A54 module. Timon is working on tuning the SLAC device to get it up and taking data.

11:55 Lost beam for a couple of minutes but it came back up when Carsten called.

12:52 MC called to say one of the klystrons in the injector is missing triggers so the beam might be coming in and out.

12:57 Stopped run 543 to change HV back to 100V.

13:40 Timon finished remotely tuning the SLAC module A97 but when a run was started there were no hits seen by the module and there were some infrequent readout errors (received data not valid). Timon is attempting to fix the problem but for now only the LBL module A54 is being run.

Access to remove the A97 to leave A54 by itself and changed to large tilt angle ~ -60o (had to rotate clockwise when looking from above, since mount frame would be in the beam line before sensor if rotated the other way).  Made choice of 60 degrees based on expectation of long cluster of 8x25um pixels for 100um thick sensor. More steep tilt would need to remount RD53a to the opposite side of frame to avoid frame in the beam line.  Couldn't mount 2nd sensor - missing screws to slide 2nd frame into beam line. One device at a time is better anyway to reduce the large multiple scattering at this steep angle. For this tile angle, hits will traverse pixel sideways with a path length of 25um/cos(60deg) ~29um.  Expected signal  peak for such short length is <2000e and really needs pixel threshold to be 1000e or less. Our pixel threshold is tuned for 1000e.        

The resulting ~4-5 pixel length of "clusters"  might be due to the artificial local XY coordinates are define against the 50x50 chip channel grid to achieve 25 micron pitch with staggered bumps ?    


Run_544 (13 deg)

Run_565 (62 deg)


17:50  Took a short run with cosmic telescope joining the main partition (so that trigger is driven by Ch-10 with NIM Quad discriminator cleanup). The FE-I4 data had no no double timing peak here in comparison to the standalone cosmicGui run just taken with trigger Ch-10 (TTL). This seems to indicate the standalone Ch-11 cosmic triggers was causing the double time peaks. The RTM ports were meant to be NIM so we were triggering on -ve undershoort after ripples of the main TTL trigger pulse ?    


RunStartEnd time#eventsRD53a DUTParticle rateComments
54111:4812:2210kC280hit/planeHV: 60 Lost beam for a couple minutes at beginning of run.
54212:2312:5610kC340hit/planeHV: 60
54413:0113:3410kC HV changed back to 100V for this and rest of runs
54613:4013:451.5kC, D Timon tuned SLAC module(D) but when the run was started no hits were seen and some sparse readout errors showed up.
55615:4516:199.9kC575hit/planeChanged angle to 60 degrees. SLAC module not connected; can't mount due to short screws. Forgot to reset histograms before start; rest around 800-900 events.
55816:3217:059.9kC20hits/planeAsked for smaller number of particles; aiming for MIMOSA->30 particles and RD53A->20.
56217:5018:032kC Added cosmic telescope to eudaq. Stopped because RCE producer died.
56819:5720:3010KC Asked for lower rate (5 particles/bcc) for tracking reconstructions. 2hits/plane for RD53A. 7hits/plane for MIMOSA.
57021:0421:3810KC Last run of the night.


21:40. We're done! Called to stop beam for entry. Going to pack up.



  • No labels